An astronaut drops a hammer from 2.0 meters above the surface of the Moon. If the acceleration due to gravity on the Moon is 1.62 meters per second2, how long will it take for the hammer to fall to the Moon's surface?
1
Expert's answer
2020-01-09T10:37:35-0500

We have:


"h=0.5g_Mt^2"

The time of flight:


"t=\sqrt{\frac{2h}{g_M}}=\sqrt{\frac{2(2)}{1.62}}=1.57\ s"